# Synthesizing data products, mathematical models, and observations for lake temperature forecasting

### v2

@@ 1 Introduction
Process-based models are critical for understanding, managing, and forecasting environmental phenomena and are widely used in ecological research (Cuddington et al., 2013). They are crucial for forecasting lake water temperature profiles (e.g., at varying depth) that influence decisions that impact the health of aquatic ecosystems (Carey et al., 2022). Water quality and the health of aquatic organisms are strongly associated with water temperature (Carey et al., 2022), which influences the growth and distribution of aquatic flora and fauna (Wetzel, 2001). Warmer water can accelerate the growth of harmful phytoplankton, threatening aquatic and human health (Carey et al., 2012). Lake temperature forecasts that extend at least 30 days into the future, include uncertainty quantification (UQ), and are updated frequently are crucial to adequately managing drinking water resources, mitigating water quality degradation, and planning for the health of lake ecosystems (Thomas et al., 2023; Carey et al., 2022). Here, we focus on forecasting lake temperatures at ten depths at Falling Creek Reservoir (FCR), a small reservoir in Vinton, Virginia, USA (Figure 1 left).
@@
There are several approaches to forecasting lake temperatures (Lofton et al., 2023), the simplest of which are climatological. These are based on historical data, capturing broader trends, and are accurate at long (15+ day) horizons (Thomas et al., 2023). Such forecasts for FCR can be constructed using data collected by in situ water temperature sensors (denoted by the red dot in Figure 1 left). Climatological forecasts do not vary substantively year upon year, their predictive limits do not vary with the forecast horizon, and they lack sharpness (Gneiting, Balabdaoui and Raftery, 2007), reducing their value for near-term water quality management.
Lake ecosystem models are designed to simulate heat transfer within aquatic ecosystems (Mooij et al., 2010) and can encompass a wide range of processes, including hydrodynamics (Hipsey et al., 2019), water quality (Hipsey et al., 2012; Cole and Wells, 2021), nutrient cycling (Janse, Aldenberg and Kramer, 1992), and ecological interactions (Leon et al., 2006). One such model, the one-dimensional (1D) General Lake Model (GLM; Hipsey et al., 2019), [^1] Apologies for the acronym; here “GLM” refers to the General Lake Model, not a generalized linear model. can simulate thermal dynamics in a vertical column of a waterbody. GLM takes into account the physical characteristics of the lake, such as its depth, surface area, and geometry, water inputs from surrounding areas, and external factors (e.g., meteorology) to simulate water movement, mixing, and circulation patterns within the water column (Hipsey et al., 2019). GLM has been used to predict long-term lake ecosystem responses to climate change with climate projection driver data (Christianson, Johnson and Hooten, 2020; Gal et al., 2020; Fenocchi et al., 2018). It has also been used to study the effects of climate change on historical lake dynamics (Ladwig et al., 2018; Shikhani et al., 2022) as well as to study lake hydrodynamics (Huang et al., 2017; Bueche, Hamilton and Vetter, 2017) and water quality indicators (Ladwig et al., 2021; Ward et al., 2020).
@@
However, the GLM has characteristics that make it challenging to use for near-term (1–30 days) forecasts of lake temperatures. Since GLM is a deterministic model, UQ associated with forecasts can only be derived if it is driven with an ensemble of model inputs, such as weather forecasts or model parameters. Importantly, like all process-based models, GLM is a simplified version of reality that does not perfectly represent processes that govern energy transfer within a water column, let alone in an entire (three-dimensional) lake. Any systematic biases, or other errors that result from such oversimplifications, need to be accounted for.
@@
Figure 1: Left: Falling Creek Reservoir (FCR; 37.30N, 79.84W). Red dot denotes sensor location. Lake depth (m) is shown by the graduated color scale. Right: Example GLM simulations (thick yellow) and observations (thick black) for 2021 at 6m; means and 90% PIs of a GLM surrogate (thinner blue); a bias-corrected (thinner pink).
@@
Observations provide a critical foundation for assessing the accuracy of GLM. At FCR, water temperature sensors installed at 10 depths (surface to 9m depth by 1m increments) provide rich and continuous observational data record starting in July 2018 (red dot in Figure 1 left). One can get a sense of GLM performance for the FCR study region by comparing GLM output to sensor data. Figure 1 (right) shows an example of GLM simulations (thick yellow line) and sensor observations (thick black line) at 6m depth as well as a surrogate model for GLM in thinner blue lines (discussed momentarily). These GLM simulations exhibit a pronounced warm bias between April and October with respect to observations.
@@
Figure 3: NOAA-GLM lake temperature forecasts (C) for FCR starting on reference dates A) 2020-10-30 and B) 2021-01-01. Colors denote trajectories of individual ensemble members. Panel numbers denote depth (m).
A more robust alternative to generating lake temperature forecasts involves building a probabilistic representation of NOAA-GLM, which could subsequently improve prediction accuracy and UQ, while also denoising the 31 ensemble forecast trajectories. This approach could accommodate functional forms of bias and correction. Modeling bias is also advantageous because it allows the bias to be separated from observational noise (Liu et al., 2009). Observe in Figure 1 how one type of statistical model (blue lines) adequately smooths GLM simulations while providing appropriate UQ. We propose to extend this idea from single to ensemble GLM forecast trajectories. Figure 1 (right) also shows the mean and 90% prediction intervals resulting from a bias-corrected model (pink lines). In contrast to the biased model (thin blue lines) fitted to GLM simulations, the mean of the bias-corrected model provides a very good approximation to observations, and UQ is reliable.
One way to statistically model NOAA-GLM forecasts is through the use of a computer model surrogate. A surrogate emulates the behavior of a computer model such as GLM but does so without the model’s physical equations (Santner et al., 2018). Gaussian process (GP) surrogates are the canonical choice as they have many desirable properties, including flexibility, accuracy, and reliable UQ (Gramacy, 2020; Williams and Rasmussen, 2006). Bias can also be modeled with a GP (Kennedy and O’Hagan, 2001; Liu et al., 2009). One major challenge to using a GP surrogate is the decomposition bottleneck of multivariate normal (MVN) covariance matrices. In GP fitting/prediction, these matrices are sized commensurately with the square of the training data of size $n$, incurring a decomposition cost that is cubic in $n$ flops. In our application, $n$ is very large. Each of the 31 NOAA ensemble members results in 31 NOAA-GLM forecasts for 30 horizons and 10 lake depths, producing a total dataset size of: 900 reference dates $\times$ 30 horizons $\times$ 10 depths $\times$ 31 ensemble members $\rightarrow n=8{,}370{,}000$, a formidable data size for any model-fitting enterprise, and especially so for GP surrogates.
@@ 2.1 Gaussian process surrogate modeling
$q(\mathbf{x}_{\_}i,\mathbf{x}_{\_}j)=\sum_{\_}{\ell=1}^{p}\left(\frac{||x_{\_}{i,\ell}-x_{\_}{j,\ell}||^{2}}{\gamma_{\_}{\ell}}\right)^{1/2}.$ (2)
@@
In what follows we express the setup in Eq. (1) concisely as $\mathcal{GP}(D_{\_}{n_{\_}M})$, given a choice of $k(\cdot)$, distance $q(\cdot)$ and hyperparameters $\boldsymbol{\theta}=(\boldsymbol{\gamma},g,\tau^{2})$. The $\boldsymbol{\gamma}=(\gamma_{\_}1,\gamma_{\_}2,\dots\gamma_{\_}{d})$ are so-called lengthscale or range parameters, where $1/\gamma_{\_}{\ell}$ determines the “relevance” of the $x_{\_}{\ell}$ input (e.g., Liu et al., 2019). The overall amplitude of the response is determined by scale parameter $\tau^{2}$, and the nugget $g$ operates as a dial partitioning that amplitude between signal $(\tau^{2})$ and noise $(\tau^{2}g)$. Observe that $g$ is only augmenting the diagonal of $\Sigma$, implementing an independent and constant (i.e., homoskedastic) noise component, an aspect which we relax in Section 3.
@@
Note that $\mathbb{V}\mathrm{ar}\{\boldsymbol{\mu}_{\_}{n_{\_}M}\}<\Sigma_{\_}{n_{\_}M}(\mathcal{X})$, uniformly over all entries of those matrices. [^2] Note also that a non-zero nugget $g$ is present elsewhere in Eq. (5). Moreover $\mathbb{V}\mathrm{ar}\{\boldsymbol{\mu}_{\_}{n_{\_}M}\}\rightarrow\boldsymbol{0}$ as $n_{\_}M\rightarrow\infty$ as long as $\mathbf{X}_{\_}{n_{\_}M}$ fills out the input space. In other words, we can learn to emulate $f$ with the surrogate $\hat{f}$ perfectly with enough data. We compare and contrast of PIs and CIs in Figure 5, alongside our heteroskedastic models in Section 3.2.
@@ 2.2 Gaussian process surrogates for lake temperatures
To illustrate, consider the following. Take $f/M$ as GLM and run it for each day of 2021, driven by observed environmental variables in that year, and extract temperatures for ten depths: surface/0 to 9m. Using the notation established above, $\mathbf{Y}_{\_}{n_{\_}M}$ is comprised of $n_{\_}M=366\times 10=3660$ simulation outputs, and $\mathbf{X}_{\_}{n_{\_}M}$ has $p=2$ columns indexing time $t$ in Julian day, and depth $d$, in meters, respectively. Figure 1 (right) shows a “slice” of some of the elements of this data ($D_{\_}{n_{\_}M}$) as a yellow line where $\mathbf{X}_{\_}{n_{\_}M}[\cdot,2]=6$ meters. (The full $D_{\_}{n_{\_}M}$ is ten times bigger.)
@@
Figure 4 (left) shows the resulting surrogate mean $\mu_{\_}{n_{\_}M}(\mathcal{X})$ over depth (m) and day of year. This corresponds to the solid blue lines in the right panel of Figure 1, which provides a slice at 6m ($\mathcal{X}[\cdot,2]=6$) along with 90% PIs. Observe in Figure 1 (right) how the surrogate smooths over GLM simulations and that intervals have appropriate coverage.
Predictions from the surrogate $\hat{f}(\mathcal{X})$ in Figures 1 (right) and 4 (left) are not “forecasts” because they do not correspond to the future. These predictions are derived from known/measured environmental variables, that, for forecasting, would be unknown/require additional modeling. That is the subject of Section 3. However, such a setup could easily be extended to serve as a surrogate for a “climatology”, if data were GLM simulations driven by observed environmental variables over a period of several years. The data setup would be the same, with two inputs and one output, but with much larger $n_{\_}M$. We do not illustrate that here, but we do compare to a similar model later in our empirical work in Section 5.
Importantly, $\hat{f}(\mathcal{X})$ in Figures 14 allow us to abstract raw GLM simulations by smoothing over jitters that come from noisy or chaotic measurements of environmental variables. Like any statistical model, the surrogate helps us separate signal from noise. Once we have the signal, we can compare it to something else, like actual physical/sensor measurements of lake temperature (black line in Figure 1). From Figure 1 we see that GLM is biased from April to November. In fact, narrow PIs here provide evidence of a strong, statistically significant bias since the black lines go outside the blue-dashed PIs in early 2021. This warm bias would have been hard to separate from noise without the surrogate – in particular without good UQ. Meanwhile, in late 2021, the situation is different: GLM is likely unbiased; any differences with sensor measurements are probably due to noise.
@@
We may then form a data set of observed discrepancies between model and field data $D^{b}_{\_}{n_{\_}F}=(\mathbf{X}_{\_}{n_{\_}F},\mathbf{Y}^{b}_{\_}{n_{\_}F})$. Then, by regarding bias, or bias correction, $b(\cdot)$ as an unknown function like $f(\cdot)$, we may model it as $\mathcal{GP}(D^{b}_{\_}{n_{\_}F})$ and infer its hyperparameters, etc. Readers familiar with the computer modeling literature will recognize a similarity between this setup and the so-called Kennedy and O’Hagan framework (KOH; Kennedy and O’Hagan, 2001) where the two processes: computer model $M$ and field data $F$ are meta-modeled jointly, without directly calculating discrepancies. Our setup here is more modular (Liu et al., 2009), and thus more amenable to up-scaling as described in Section 3. Also we are not learning any so-called calibration parameters, although we may explore that in future work (Appendix E), so there is no potential for confounding (e.g., Brynjarsdóttir and O’Hagan, 2014).
@@
Now consider Figure 1 (right), which shows the bias correction, in pink, formed as the sum of two GP fits: GLM surrogate $\hat{f}$ and bias correction $\hat{b}$. One may think of this as (a slice through) the sum of the first two panels in Figure 4, but that would only account for means. A more accurate depiction would involve 2d versions of the slices through the full distribution(s) summarized by the blue lines ($\hat{f}$) in Figure 1 (right) and for the bias correction $\hat{b}$ in the right panel of Figure 4. Mathematically, since $\hat{f}$ and $\hat{b}$ are both MVNs we can follow formulas for sums of MVNs which involve summing the means $(\mu_{\_}{n_{\_}M}(\mathcal{X})+\mu^{b}_{\_}{n_{\_}F}(\mathcal{X})$) and covariances ($\mathbb{V}\mathrm{ar}\{\mu_{\_}{n_{\_}M}(\mathcal{X})\}+\Sigma^{b}_{\_}{n_{\_}F}(\mathcal{X})$) and subtracting off cross-covariances. The pink lines shown in Figure 1 omit subtracted cross-covariances, somewhat overestimating uncertainty. The need to subtract cross-covarainces is due to a negative correlation between the GLM surrogate and bias GP (e.g. as one increases, the other would likely decrease) (Kennedy and O’Hagan, 2001; Higdon et al., 2004). However, we cannot access cross-covariances due the R implementation that we use for fitting highly efficient GPs throughout this work (see Section 3.3). Consequently, the total variance of the bias-corrected surrogate is over-inflated slightly. However, we do not believe the over-inflation of this variance negatively affects UQ, as we would prefer conservative UQ over over-confidence of forecasts.
It is important to note that we are not using $\Sigma_{\_}{n_{\_}M}(\mathcal{X})$, the full predictive variance for $Y(\mathcal{X}\mid D_{\_}{n_{\_}M})$ from Eq. (4) for the bias corrected (co-) variance. Instead we are using the nugget-free one (5) because only the estimated GLM surrogate mean $\boldsymbol{\mu}_{\_}{n_{\_}M}$ was used to define the observed discrepancies, not actual GLM simulations. We shall delve into this further in Section 3.2. For more details, see Gramacy (2020, Sections 5.3.2 & 8.1.3). Observe from Figure 1 (right) that the resulting predictions and PIs in pink are accurate and provide good coverage for the field data. This is perhaps not surprising as we are predicting in-sample, and the GP (or two GPs in fact) provide in a highly flexible non-linear regression. It is their performance out of sample, for future lake temperatures, which is our focus.

### v3

@@ 1 Introduction
Process-based models are critical for understanding, managing, and forecasting environmental phenomena and are widely used in ecological research. They are crucial for forecasting lake water temperature profiles (e.g., at varying depth) that influence decisions that impact the health of aquatic ecosystems (Carey et al., 2022). Water quality and the health of aquatic organisms are strongly associated with water temperature (Carey et al., 2022), which influences the growth and distribution of aquatic flora and fauna (Wetzel, 2001). Warmer water can accelerate the growth of harmful phytoplankton, threatening aquatic and human health (Carey et al., 2012). Lake temperature forecasts that extend at least 30 days into the future, include uncertainty quantification (UQ), and are updated frequently are crucial to adequately managing drinking water resources, mitigating water quality degradation, and planning for the health of lake ecosystems (Thomas et al., 2023; Carey et al., 2022). Here, we focus on forecasting lake temperatures at ten depths at Falling Creek Reservoir (FCR), a small reservoir in Vinton, Virginia, USA (Figure 1 left).
@@
There are several approaches to forecasting lake temperatures (Lofton et al., 2023), the simplest of which are climatological. These are based on historical data, capturing broader trends, and are accurate at long (15+ day) horizons (Thomas et al., 2023). Such forecasts for FCR can be constructed using data collected by in situ water temperature sensors (denoted by the red dot in Figure 1 left). Climatological forecasts do not vary substantively year upon year, their predictive limits do not vary with the forecast horizon, and they lack sharpness reducing their value for near-term water quality management.
@@
Lake ecosystem models are designed to simulate heat transfer within aquatic ecosystems and can encompass a wide range of processes, including hydrodynamics, water quality, nutrient cycling, and ecological interactions. One such model, the one-dimensional (1D) General Lake Model (GLM; Hipsey et al., 2019), [^1] Apologies for the acronym; here “GLM” refers to the General Lake Model, not a generalized linear model. can simulate thermal dynamics in a vertical column of a waterbody. GLM takes into account the physical characteristics of the lake, such as its depth, surface area, and geometry, water inputs from surrounding areas, and external factors (e.g., meteorology) to simulate water movement, mixing, and circulation patterns within the water column (Hipsey et al., 2019). GLM has been used in many contexts, which we do not review here. However, the GLM has characteristics that make it challenging to use for near-term (1–30 days) forecasts of lake temperatures. Since GLM is a deterministic model, UQ associated with forecasts can only be derived if it is driven with an ensemble of model inputs, such as weather forecasts or model parameters. Importantly, like all process-based models, GLM is a simplified version of reality that does not perfectly represent processes that govern energy transfer within a water column. Any systematic biases, or other errors that result from such oversimplifications, need to be accounted for.
@@
Figure 1: Left: Falling Creek Reservoir (FCR; 37.30N, 79.84W). Right: Example GLM simulations (lighter thick lines) and observations (thick dark lines) for 2021 at 6m; means and 90% PIs of a GLM surrogate (thin lines); a bias-corrected (light, thin lines).
@@
Observations provide a critical foundation for assessing the accuracy of GLM. At FCR, water temperature sensors installed at 10 depths (surface to 9m depth by 1m increments) provide a rich and continuous observational data record starting in July 2018 (dot in Figure 1 left). One can get a sense of GLM performance for the FCR study region by comparing GLM output to sensor data. Figure 1 (right) shows an example of GLM simulations (thick, light line) and sensor observations (thick, dark line) at 6m depth as well as a surrogate model for GLM in thinner lines (discussed momentarily). These GLM simulations exhibit a pronounced warm bias between April and October with respect to observations.
@@
Figure 3: NOAA-GLM lake temperature forecasts (C) for FCR starting on reference dates A) 2020-10-30 and B) 2021-01-01. Different shades denote trajectories of ensemble members. Panel numbers denote depth (m).
A more robust alternative to generating lake temperature forecasts involves building a probabilistic representation of NOAA-GLM, which could subsequently improve prediction accuracy and UQ, while also denoising the 31 ensemble forecast trajectories. This approach could accommodate functional forms of bias and correction. Modeling bias is also advantageous because it allows the bias to be separated from observational noise. Observe in Figure 1 how one type of statistical model (the surrogate) adequately smooths GLM simulations while providing appropriate UQ. We propose to extend this idea from single to ensemble GLM forecast trajectories. Figure 1 (right) also shows the mean and 90% prediction intervals resulting from a bias-corrected model. In contrast to the biased surrogate model fitted to GLM simulations, the mean of the bias-corrected model provides a very good approximation to observations, and UQ is reliable.
One way to statistically model NOAA-GLM forecasts is through the use of a computer model surrogate. A surrogate emulates the behavior of a computer model such as GLM but does so without the model’s physical equations (Santner et al., 2018). We propose using a surrogate to model NOAA-GLM lake temperature forecasts. Gaussian process (GP) surrogates are the canonical choice as they have many desirable properties, including flexibility, accuracy, and reliable UQ (Gramacy, 2020; Williams and Rasmussen, 2006). Bias can also be modeled with a GP (Kennedy and O’Hagan, 2001). One major challenge to using a GP surrogate is the decomposition bottleneck of multivariate normal (MVN) covariance matrices. In GP fitting/prediction, these matrices are sized commensurately with the square of the training data of size $n$, incurring a decomposition cost that is cubic in $n$ flops. In our application, $n$ is very large. Each of the 31 NOAA ensemble members results in 31 NOAA-GLM forecasts for 30 horizons and 10 lake depths, producing a total dataset size of: 900 reference dates $\times$ 30 horizons $\times$ 10 depths $\times$ 31 ensemble members $\rightarrow n=8{,}370{,}000$, a formidable data size for any model-fitting enterprise, especially for GP surrogates.
@@ 2.1 Gaussian process surrogate modeling
$q(\mathbf{x}_{\_}i,\mathbf{x}_{\_}j)=\left(\sum_{\_}{\ell=1}^{p}\frac{||x_{\_}{i,\ell}-x_{\_}{j,\ell}||^{2}}{\gamma_{\_}{\ell}}\right)^{1/2}.$ (2)
@@
In what follows we express the setup in Eq. (1) concisely as $\mathcal{GP}(D_{\_}{n_{\_}M})$, given a choice of $k(\cdot)$, distance $q(\cdot)$ and hyperparameters $\boldsymbol{\theta}=(\boldsymbol{\gamma},g,\tau^{2})$. The $\boldsymbol{\gamma}=(\gamma_{\_}1,\gamma_{\_}2,\dots\gamma_{\_}{d})$ are so-called lengthscale or range parameters, where $1/\gamma_{\_}{\ell}$ determines the “relevance” of the $x_{\_}{\ell}$ The overall amplitude of the response is determined by scale parameter $\tau^{2}$, and the nugget $g$ operates as a dial partitioning that amplitude between signal $(\tau^{2})$ and noise $(\tau^{2}g)$. Observe that $g$ is only augmenting the diagonal of $\Sigma$, implementing an independent and constant (i.e., homoskedastic) noise component, an aspect that we relax in Section 3.
@@
Note that $\mathbb{V}\mathrm{ar}\{\boldsymbol{\mu}_{\_}{n_{\_}M}\}<\Sigma_{\_}{n_{\_}M}(\mathcal{X})$, uniformly over all entries of those matrices. [^2] Note also that a non-zero nugget $g$ is present elsewhere in Eq. (5). Moreover $\mathbb{V}\mathrm{ar}\{\boldsymbol{\mu}_{\_}{n_{\_}M}\}\rightarrow\boldsymbol{0}$ as $n_{\_}M\rightarrow\infty$ as long as $\mathbf{X}_{\_}{n_{\_}M}$ fills out the input space. In other words, we can learn to emulate $f$ with the surrogate $\hat{f}$ perfectly with enough data. We compare and contrast PIs and CIs in Figure 5, alongside our heteroskedastic models in Section 3.2.
@@ 2.2 Gaussian process surrogates for lake temperatures
To illustrate, consider the following. Take $f/M$ as GLM and run it for each day of 2021, driven by observed environmental variables in that year, and extract temperatures for ten depths: surface/0 to 9m. Using the notation established above, $\mathbf{Y}_{\_}{n_{\_}M}$ is comprised of $n_{\_}M=366\times 10=3660$ simulation outputs, and $\mathbf{X}_{\_}{n_{\_}M}$ has $p=2$ columns indexing time $t$ in Julian day, and depth $d$, in meters, respectively. Figure 1 (right) shows a “slice” of some of the elements of this data ($D_{\_}{n_{\_}M}$) where $\mathbf{X}_{\_}{n_{\_}M}[\cdot,2]=6$ meters. (The full $D_{\_}{n_{\_}M}$ is ten times bigger.)
@@
Figure 4 (left) shows the resulting surrogate mean $\mu_{\_}{n_{\_}M}(\mathcal{X})$ over depth (m) and day of year. This corresponds to the dark lines in the right panel of Figure 1, which provides a slice at 6m ($\mathcal{X}[\cdot,2]=6$) along with 90% PIs. Observe in Figure 1 (right) how the surrogate smooths over GLM simulations and that intervals have appropriate coverage.
Predictions from the surrogate $\hat{f}(\mathcal{X})$ in Figures 1 (right) and 4 (left) are not “forecasts” because they do not correspond to the future. These predictions are derived from known or measured environmental variables, that, for forecasting, would be unknown or require additional modeling. That is the subject of Section 3.
Importantly, $\hat{f}(\mathcal{X})$ in Figures 14 allow us to abstract raw GLM simulations by smoothing over noisy or chaotic measurements of environmental variables. Like any statistical model, the surrogate helps us separate signal from noise. Once we have the signal, we can compare it to something else, like actual physical/sensor measurements of lake temperature (black line in Figure 1). From Figure 1 we see that GLM is biased from April to November. In fact, narrow PIs here provide evidence of a strong, statistically significant bias since the black lines go outside the blue-dashed PIs in early 2021. This warm bias would have been hard to separate from noise without the surrogate – in particular without good UQ. Meanwhile, in late 2021, the situation is different: GLM is likely unbiased; any differences with sensor measurements are probably due to noise.
@@
We may then form a data set of observed discrepancies between model and field data $D^{b}_{\_}{n_{\_}F}=(\mathbf{X}_{\_}{n_{\_}F},\mathbf{Y}^{b}_{\_}{n_{\_}F})$. Then, by regarding bias, or bias correction, $b(\cdot)$ as an unknown function like $f(\cdot)$, we may model it as $\mathcal{GP}(D^{b}_{\_}{n_{\_}F})$ and infer its hyperparameters, etc. Readers familiar with the computer modeling literature will recognize a similarity between this setup and the so-called Kennedy and O’Hagan framework (KOH; Kennedy and O’Hagan, 2001) where the two processes: computer model $M$ and field data $F$ are meta-modeled jointly, without directly calculating discrepancies. Our setup here is more modular, and thus more amenable to up-scaling as described in Section 3.
@@
Now consider Figure 1 (right), which shows the bias correction formed as the sum of two GP fits: GLM surrogate $\hat{f}$ and bias correction $\hat{b}$. One may think of this as (a slice through) the sum of the first two panels in Figure 4, but that would only account for means. A more accurate depiction would involve 2d versions of the slices through the full distribution(s) summarized by $\hat{f}$ in Figure 1 (right) and for the bias correction $\hat{b}$ in the right panel of Figure 4. Mathematically, since $\hat{f}$ and $\hat{b}$ are both MVNs we can follow formulas for sums of MVNs which involve summing the means $(\mu_{\_}{n_{\_}M}(\mathcal{X})+\mu^{b}_{\_}{n_{\_}F}(\mathcal{X})$) and covariances ($\mathbb{V}\mathrm{ar}\{\mu_{\_}{n_{\_}M}(\mathcal{X})\}+\Sigma^{b}_{\_}{n_{\_}F}(\mathcal{X})$) and subtracting off cross-covariances. The bias-corrected surrogate in Figure 1 omits subtracted cross-covariances due to a negative correlation between the GLM surrogate and bias GP, somewhat overestimating uncertainty. However, we do not believe the over-inflation of the total variance negatively affects UQ, as we would prefer conservative UQ over over-confidence of forecasts.
It is important to note that we are not using $\Sigma_{\_}{n_{\_}M}(\mathcal{X})$, the full predictive variance for $Y(\mathcal{X}\mid D_{\_}{n_{\_}M})$ from Eq. (4) for the bias corrected (co-) variance. Instead we are using the nugget-free one (5) because only the estimated GLM surrogate mean $\boldsymbol{\mu}_{\_}{n_{\_}M}$ was used to define the observed discrepancies, not actual GLM simulations. We shall delve into this further in Section 3.2. For more details, see Gramacy (2020, Sections 5.3.2 & 8.1.3). Observe from Figure 1 (right) that the resulting predictions and PIs are accurate and provide good coverage for the field data. This is perhaps not surprising as we are predicting in-sample, and the GP (or two GPs in fact) provide a highly flexible non-linear regression. It is their performance out of sample, for future lake temperatures, which is our focus.